OpenStudy (anonymous):

Evaluate the function rule to find the range of each function for the domain {-3, 0, 5}. f(x) = x² – 5

OpenStudy (anonymous):

plug in each of those numbers in the domain into the function to get your range....

OpenStudy (anonymous):

for example... x=-3: \[\huge f(-3)=(-3)^2-5=9-5=4 \] so 4 is a part of the range.

OpenStudy (anonymous):

do that with the other two numbers in the domain....

OpenStudy (anonymous):

4 , 5 , 10 ?

OpenStudy (anonymous):

no... your second and third numbers are incorrect...